OpenStudy (anonymous):

The waitress and the hostess at a restaurant share the tips at the end of the shift. The average amount earned in tips between the two of them is $75. The waitress gets $45 and the hostess gets $30. If they continue to earn money at this rate, how many dollars will the waitress receive if they earn $350 at the end of the shift?

OpenStudy (anonymous):

is this multiple choice?

OpenStudy (anonymous):

@turtlemaroon Yes, but I want it to be explained to me, I don't want someone to just randomly choose one of the multiple choice and give me the wrong answer or something.

OpenStudy (anonymous):

What I do is figure it out myself, and see if the answer I get is one of the choices. I never randomly choose. I only answer if I know for certain. :)

OpenStudy (anonymous):

and I also type out how I found the answer here

OpenStudy (anonymous):

You don't have to tell me which one is correct if you don't want to. I just want you to explain to me how you got the answer. A.$225 B.$216 C.$210 D.$213

OpenStudy (anonymous):

mkay, thanks. Give me a minute, I'm going to try and figure it out :)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Okay, I know the answer. I figured it out in an unconventional way, though. Would you like to know how I did it?

OpenStudy (anonymous):

Yes please. ^^

OpenStudy (anonymous):

Okay, I divided 45 by 75. That is 0.6. Now, I take each of the answer choices, and divide each by 350. I find that 210/350 is also .6. If you know simple fractions (like 1/2 = 4/8), you know that 45/75 and 210/350 are equal, so C must be the answer!

OpenStudy (anonymous):

Actually I just solved it. It finally just made sense to me. I wrote a proportion for it, your answer is correct, thank you. \[\frac{ 45 }{ 75 }=\frac{ n }{ 100 } \] and then cross multiply. \[45\times350=75\times n\]\[15,750=75n\] Now divide 15,750 and 75 by 75 to isolate and determine the unknown variable. \[\frac{ 15,750 }{ 75 }=\frac{ 75n }{ 75 }\]

OpenStudy (anonymous):

@turtlemaroon, After you work that out 75 is canceled out and all you have left is the quotient from 15,750 divided by 75 which equals $210. \[\frac{ 15,750 }{ 75 } =210 --> 15,750\div75=210\] Thank you though! :)